# Evaluating End-User Device Energy Models in Sustainability Reporting of Browser-Based Web Services

### v1

@@
Abstract.
Sustainability reporting in web-based services increasingly relies on simplified energy and carbon models such as the Danish Agency of Digital Government’s Digst framework and the United Kingdom-based DIMPACT model. Although these models are widely adopted, their accuracy and precision remain underexplored. This paper presents an empirical study evaluating how well such models reflect actual energy consumption during realistic user interactions with common website categories. Energy use was measured across shopping, booking, navigation, and news services using predefined user flows executed on four laptop platforms. The results show that the commonly applied constant-power approximation ($P\times t$) can diverge substantially from measured energy, depending on website category, device type, and task characteristics. The findings demonstrate that model deviations are systematic rather than random and highlight the need for category-aware and device-reflective power parameters in reproducible sustainability reporting frameworks.
Energy Consumption, End-user device, Energy Models, Reporting
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 41st ACM/SIGAPP Symposium on Applied Computing; March 23–27, 2026; Thessaloniki, Greeceisbn: 979-X-XXXX-XXXX-X/26/03ccs: General and reference Evaluationccs: Software and its engineering Empirical software validation
1. Introduction
Digital sustainability reporting has become central to the evaluation of web-based services. Governments and organizations increasingly depend on simplified models to estimate energy use and greenhouse gas emissions. Two such models are the Danish Agency of Digital Government’s Digst framework (digst, 7) and the UK-based DIMPACT framework (dimpact, 8, 26) based on (Preist2019, 24). However, despite their growing role, the empirical validation of these models is limited. Without systematic assessment, it remains unclear how accurately they capture the energy demands of real-world web interactions.
This paper addresses this gap by presenting an experiment that directly compares measured energy consumption against model-based estimates. Following the recommendation by Guldner et al. (Guldner2024, 15) to evaluate software energy efficiency under representative user scenarios, and inspired by the online-media categories proposed by Schien et al. (Schien2013, 27)—who modeled energy use across text-based, interactive, and streaming web content—we focus on non-streaming web services (shopping, booking, navigation, and news). In doing so, we complement prior work that has predominantly investigated streaming services. While categories such as streaming and generative AI are highly relevant to digital sustainability, their energy use depends on complex scaling parameters (e.g., video quality (Suski2020, 28) or token length) that require separate, parameter-specific studies. We therefore delimit our study to web services with more stable and comparable interaction patterns.
@@ 1. Introduction
    How accurately do current reporting models estimate the energy consumption of end-user devices during realistic web-service interactions?
@@
    Which factors—such as device type, website category, or interaction duration—explain systematic deviations between modeled and measured energy use?
Our contributions are threefold: (1) a reproducible experimental methodology for measuring the energy use of web-service interactions, (2) empirical evidence on the accuracy and precision of Digst- and DIMPACT-style energy estimates, and (3) insights for improving fairness and transparency in digital sustainability reporting.
2. Background and Related Work
@@
This section situates the study within existing research on the estimation and reporting of ICT-related energy use. We first review sector-level and national models that underpin current digital sustainability frameworks, focusing on their assumptions about end-user devices. We then outline software- and measurement-level approaches that aim to quantify energy use more precisely at the program or interaction level. Together, these strands of work frame the methodological gap that this paper addresses: linking reproducible, software-level energy measurements to the simplified models used in large-scale sustainability reporting.
2.1. Sector-Level Models
@@
Global and sector-level ICT footprint models establish the broader context for digital sustainability research. Early studies, e.g., (Andrae2015, 1, 21, 2), raised awareness of the growing energy challenges, while Freitag et al. (Freitag2021, 11) analyzed the methodological assumptions that explain divergent estimates of Internet energy intensity. Later studies, such as Malmodin et al. (malmodin2024, 20), provide comprehensive assessments of the ICT sector’s total energy use and greenhouse gas emissions, emphasizing harmonized system boundaries across data centers, networks, and end-user devices.
@@
These models typically employ top-down allocation and averaged intensity factors to distribute total sector energy across subdomains. As a result, they inherently contain uncertainties, and Malmodin et al. specifically acknowledge that “the number of user devices in operation and their average electricity consumption remain a more uncertain part in line with previous studies” (malmodin2024, 20). This uncertainty is echoed by Furberg et al. (Furberg2026, 12), who note that “to estimate the electricity use of devices, data on actual device usage patterns (e.g., operational service time per year) are required. In practice, studies on the ICT sector’s direct climate impact usually do not collect device usage data but rely on secondary data from various sources.”e.g., averages over multiple sources (malmodin2024, 20).
Although these studies establish the overall scale of ICT’s environmental impact, their level of aggregation prevents validation of individual services or user interactions. Consequently, sector models cannot determine whether specific web services, user actions, or design choices meaningfully reduce energy demand. These limitations motivate the development of service-level models for digital sustainability reporting, which attempt to translate aggregate insights into interpretable metrics for individual services.
2.2. Service-Level Models
@@
Service-level models aim to make the environmental assessment of digital services practical and comparable. Frameworks such as the Danish Agency of Digital Government [^1] https://en.digst.dk/’s Digst model (digst, 7) and the United Kingdom-based DIMPACT tool (dimpact, 8, 26) operationalize simplified estimation methods that express energy use as the product of assumed device power and activity duration ($P\times t$). These models allocate a share of data-center and network energy to each service and are increasingly used for organizational sustainability reporting. The approach aligns with academic efforts such as Preist et al. (Preist2019, 24), who analyzed how interface design in YouTube affects energy demand, and with the Green Software Foundation’s Software Carbon Intensity (SCI) specification, which formalizes functional units and allocation principles for shared infrastructure.
@@
Although these frameworks facilitate transparency and comparability, they rely on several simplifying assumptions: constant device power, uniform user behavior, and linear relationships between time, data volume, and energy use.
@@
While these models target sustainability reporting, a parallel line of work addresses developer-facing concerns by improving software energy efficiency and promoting reproducible energy-measurement practices.
2.3. Software-Level Efficiency and Measurement Frameworks
Research on software-level energy efficiency has produced a range of tools and methodologies for analyzing, predicting, and improving the energy behavior of programs. Early frameworks such as EACOF (Field2014, 10) enabled applications to access power data at runtime, fostering awareness of energy use during development. Static and data-dependent analyses by Grech et al. (Grech2015, 14) and Pallister et al. (Pallister2017, 22) estimate energy consumption directly from program structure or input characteristics, while comparative studies such as Georgiou et al. (Georgiou2017, 13) and Pereira et al. (Pereira2021, 23) quantify how programming language choice influences energy efficiency. Complementary approaches examine reusable solutions for efficient design—such as energy-related software patterns (Rani2024, 25, 9)—and energy-aware debugging and testing frameworks for mobile and desktop applications.
@@ 2.3. Software-Level Efficiency and Measurement Frameworks
More recently, Guldner et al. (Guldner2024, 15) proposed the Green Software Measurement Model, a standardized process for measuring software resource efficiency under representative user scenarios. Their rationale is that meaningful and reproducible results require measurements grounded in realistic user interactions rather than synthetic benchmarks. This principle directly aligns with our methodological choice to evaluate energy consumption through end-to-end user flows across web services.
While these works provide methodological foundations for measuring and improving software efficiency, their focus remains on optimizing or characterizing software behavior in isolation. Our study builds upon their measurement philosophy but applies it to a different objective: validating the accuracy of service-level reporting models on real-world services.
2.4. Web-Specific Energy Studies
Energy consumption in web-based services has been investigated from multiple perspectives, ranging from modeling of user-device interactions to experimental measurement of browser and video workloads. Schien et al. (Schien2013, 27) provided one of the earliest models of online media consumption, distinguishing between text-based, interactive, and streaming content across devices such as desktops, laptops, and mobile phones. Their results revealed that end-user devices, particularly laptops, dominate overall energy consumption, and that usage patterns substantially influence total impact. This modeling work established a foundation for categorizing web services by data intensity and interaction type—a principle we adopt when structuring our own experimental categories.
Subsequent studies examined specific behavioral and technical factors that shape web energy use. Suski et al. (Suski2020, 28) demonstrated that user decisions, such as video resolution and device choice, strongly affect streaming emissions. Rani et al. (Rani2024, 25) identified recurring energy-related design patterns in web applications, emphasizing that front-end implementation choices can meaningfully alter client-side power demand. Macedo et al. (Macedo2022, 3) compared WebAssembly and JavaScript, showing that runtime technology influences both execution time and energy efficiency. Together, these studies highlight the diversity of determinants—behavioral, architectural, and technological—that drive energy use in web contexts.
Recent work has advanced the methodological basis for measuring energy in multimedia and browser environments. Katsenou et al. (Katsenou2024, 18) compared hardware- and software-based power measurements in video compression, while Herglotz et al. (Herglotz2025, 16) formalized procedures for synchronized power and time measurements in video communication systems. Their controlled setups provide technical precision that complements our empirical approach. In our study, news-site user flows occasionally included embedded video playback, allowing us to relate these real-world observations to the controlled results reported by Herglotz et al. Similarly, our measured laptop data provide empirical confirmation of Schien et al.’s modeled dominance of device energy use.
@@ 2.4. Web-Specific Energy Studies
Collectively, this literature demonstrates that web-service energy use is heterogeneous and context dependent. However, few of these studies examine whether the simplified assumptions used in service-level reporting models can capture such variability. Our work addresses this gap by providing measured, category-based data to test the validity of those assumptions under realistic web interactions.
2.5. Positioning of This Work
Existing literature thus ranges from sector-wide modeling to code-level analysis, with few studies bridging measured data and reporting models. Our study occupies this intermediate layer: it empirically tests the accuracy and precision of service-level energy models (Digst and DIMPACT) under realistic user flows. Following the recommendation of Guldner et al. (Guldner2024, 15) and building on the content categories proposed by Schien et al. (Schien2013, 27), we structure our experiment around four non-streaming web-service categories—shopping, booking, navigation, and news—and later relate our results to prior findings on device-level energy variation and video playback effects. The following section details our experimental design, which applies these principles in a controlled setting to test model validity across four representative web-service categories.
3. Methodology
This section describes the experimental design, data collection, and model evaluation procedures used to address the research questions. We outline the selection of representative web services and user flows, the hardware and participant setup, and the measurement process for recording power and duration. Finally, we detail how the measured data were compared with the simplified energy models used in existing reporting frameworks to assess their accuracy and structural bias.
3.1. Websites and User Flows
We selected eight popular websites grouped into four categories:
  1. (1)
@@ 3.1. Websites and User Flows
  3. (3)
    Directions: google.com/maps, bing.com/maps; and
  4. (4)
@@
For each category, we defined realistic user flows that mimic everyday usage—searching for a product, reserving a hotel room, generating a route, or reading a news article—so that equivalent functionality was tested across services. Streaming and other parameter-dependent services were deliberately excluded to ensure comparability of interaction length and structure.
3.2. Devices
@@
Experiments were performed on four laptops representing different hardware generations and OEM configurations:
  1. Laptop 1:
    Dell Latitude 7490 (Intel 8th-gen, UHD 620),
  2. Laptop 2:
    HP EliteBook 830 G6 (Intel 8th-gen, alternate firmware),
  3. Laptop 3:
    Dell Latitude E5250 (Intel 5th-gen, compact chassis), and
  4. Laptop 4:
    Dell Latitude E5550 (Intel 5th-gen, larger chassis).
@@ 4. Laptop 4:
This selection captures both processor-generation and design-level power variation.
3.3. Participants and Experimental Design
@@
Ten participants completed the experiment; one participant (P10) contributed only two runs and was excluded from the mixed-effects analyses to ensure stable estimation of random effects. The remaining nine participants provided between eight and twenty-four valid runs each. The design crossed eight user flows with four laptops, and every flow–computer combination was executed by at least four and up to seven distinct participants, ensuring balanced coverage. To counter potential learning or fatigue effects while maintaining full coverage, the order of user flows and the allocation of laptops were randomized under constraints that ensured each flow–device combination was represented across participants. Each flow was repeated three times per device, yielding a total of $164$ valid observations.
3.4. Experimental Setup and Measurements
@@
All laptops ran identical operating-system images (Ubuntu Desktop 24.04.3 LTS) with default power settings and brightness. The same Chromium browser version was used in incognito mode, and caches were cleared between runs. Each measurement recorded device power (W) via a Siglent SPD3303X-E programmable power supply and task duration (s), allowing energy to be computed as integrated power over time. The external power-supply measurements was user-synchronized via task start and stop events to ensure precise alignment between user interaction and power logging. The setup can be seen in Figure1. Key sources of potential variance included machine architecture, user execution behavior, and learning effects, which were addressed through randomization and replication. Potential influences of these factors and the sample size are discussed in Section 7.
Figure 1. The setup for the experiment with 4 Laptops, two Siglents and two data-collecting computers with user-controlled synchronization software.
3.5. Model Evaluation
@@
To evaluate how well sector-level reporting models represent actual end-user device energy, we implemented both the Digst and DIMPACT formulations. In both frameworks, energy use is estimated as constant device power multiplied by interaction duration ($E_{\mathrm{model}}=P_{\mathrm{avg}}\times t$).
@@
Digst uses (digst, 7, Table 1 ) a composite averaging over laptops (22 W), tower desktops (87 W), all-in-one PCs (81 W), and monitors (31 W) to derive an overall mean user-PC power of approximately 68.5 W, i.e., the average power ratings reported by Urban et al. (urban2017, 29). However, since we focus on laptops, we have extracted the laptop power of 22 W. In comparison, DIMPACT have in the case of browser-based services chosen a lower constant of 15 W for web-browsing scenarios (dimpact, 8, 26), i.e., Urban’s power ratings without workload adjustment (urban2017, 29). We will use these constants (22W and 15W) to evaluating energy models in Section 5.1.
4. Reproducibility
All anonymized datasets, user-flow descriptions, and analysis scripts are available in the open repository EUD-energy-model [^2] https://github.com/SustainableSoftware/EUD-energy-model. The repository includes all data, scripts, version numbers, and documentation needed to reproduce the study and the analyses in this study.
5. Results
Our analysis examined accuracy and precision across categories, machines, and flows. Initial findings suggest that while power multiplied by time estimates are broadly correlated with measured energy, deviations occur systematically. In particular, shopping showed higher variance, while navigation tasks on maps platforms tasks produced more consistent estimates across devices, e.g., visualized in Figures 7 and  8. Machine differences (5th vs. 8th gen CPUs, OEM design) also contributed significantly to deviations, e.g., visualized in Figure 6.
5.1. Model Accuracy and Error Growth (RQ1)

### v2

# 1st Maja H. Kirkeby    2nd Timmie Lagermann

@@
Abstract
Sustainability reporting for web-based services often relies on simplified end-user energy models that assume constant laptop power during browser interactions. Energy models such as Digst and DIMPACT apply fixed power values (15–22 W), yet the validity of this approach for realistic browsing remains underexplored.
We empirically evaluate constant-power assumptions in a controlled user study where ten participants repeatedly complete eight representative user flows across shopping, booking, navigation, and news services on four laptop platforms, while device energy is measured. Typical power is 9–13 W, substantially below current reporting standards, implying systematic overestimation. Moreover, the error scales proportionally with task duration, indicating systematic bias rather than random noise.
Comparing progressively refined constant-power models, we find that category-specific parameters improve accuracy more than hardware-only parameters and approach flow-specific performance. The best fit is obtained by combining category (or flow) with hardware, while category-level models retain most of the benefit with fewer parameters, making them a practical upgrade for sustainability reporting.
I Introduction
Digital sustainability reporting has become central to the evaluation of web-based services. Governments and organizations increasingly depend on simplified models to estimate energy use and greenhouse gas emissions. Two such models are the Danish Agency of Digital Government’s Digst framework [1] and the UK-based DIMPACT framework [2, 3] based on [4]. However, despite their growing role, the empirical validation of these models is limited. Without systematic assessment, it remains unclear how accurately they capture the energy usage of real-world web interactions.
This paper addresses this gap by presenting an experiment that directly compares measured energy consumption against model-based estimates. Following the recommendation by Guldner et al. [5] to evaluate software energy efficiency under representative user scenarios, and inspired by the use of online-media categories by Schien et al. [6], we focus on primarily non-streaming web services: shopping, booking, navigation, and news. While categories such as audio/video streaming and generative AI are highly relevant to digital sustainability, their energy use depends on complex scaling parameters, e.g.,  [7, 8], that require separate, parameter-specific studies. We therefore delimit our study to web services with more stable and comparable interaction patterns.
@@ I Introduction
How accurately do constant-power assumptions in current reporting frameworks approximate measured end-user device energy use during realistic browser-based user flows?
@@
How does accuracy change when constant-power models are parameterized by device, service category, and user flow, and which factors contribute most to improved precision?
To answer these, we make the following contributions: (1) Protocol & dataset: a controlled method (and dataset) for measuring end-user device energy during realistic browser flows across services and laptops. (2) Duration-proportional bias: empirical evidence that constant-power assumptions induce systematic error that scales with interaction duration. (3) Reporting refinement: a quantified trade-off analysis for device/category/flow parameterization, identifying category-level calibration as a low-cost accuracy gain.
The paper is organized as follows. Section II reviews related work on ICT sustainability reporting and web-energy measurement. Section III describes the study design, measurement setup, and model evaluation procedure. Section IV presents results for RQ1 and RQ2, and Section V discusses threats to validity. Section VI interprets implications for sustainability reporting models, and Section VII concludes with future work.
II Background and Related Work
@@
This section situates the study within existing research on the estimation and reporting of Information and Communication Technology (ICT) related energy use. We will focus on how they deal with use-phase energy consumption for end-user devices. We review Sector-level and Service-level models for reporting, software-level energy-efficiency measurement frameworks, and the web-specific energy efficiency. We then outline software- and measurement-level approaches that aim to quantify energy use more precisely at the program or interaction level. Together, these works frame the methodological gap that this paper addresses: linking reproducible, software-level energy measurements to the simplified models used in large-scale sustainability reporting.
II-A Sector-Level Reporting Models
@@
Global and sector-level ICT footprint models establish the broader context for digital sustainability research. Early studies, e.g., [9, 10, 11], raised awareness of the growing energy challenges, while Freitag et al. [12] analyzed the methodological assumptions. Later studies, such as Malmodin et al. [13], provide comprehensive evaluation of the ICT sector’s total energy use and greenhouse gas emissions.
@@
These models typically employ top-down allocation and averaged intensity factors to distribute total sector energy across subdomains. As a result, they inherently contain uncertainties, and Malmodin et al. [13] specifically acknowledge that “the number of user devices in operation and their average electricity consumption remain a more uncertain part in line with previous studies”. This uncertainty is echoed by Furberg et al. [14], who note that “to estimate the electricity use of devices, data on actual device usage patterns (e.g., operational service time per year) are required. In practice, studies on the ICT sector’s direct climate impact usually do not collect device usage data but rely on secondary data from various sources”, e.g., averages over multiple sources [13].
In relation to our study, these studies provide an essential macro-level baseline for ICT sustainability. While uncertainty is inherent to top-down footprinting, it is notable that end-user devices’ electricity consumption and usage patterns are emphasized as a particularly uncertain component [13, 14]. These models relate to service-level models, which operationalize this approach in actionable estimates at the granularity of specific services and user interactions
II-B Service-Level Reporting Models
@@
Service-level models aim to make the environmental assessment of digital services practical and comparable. Frameworks such as the Danish Agency of Digital Government [^1] https://en.digst.dk/’s Digst model [1] and the United Kingdom-based DIMPACT tool [2, 3] operationalize simplified estimation methods that express energy use as the product of assumed constant device power and activity duration ($P\times t$).
These models allocate a share of data center and network energy to each service and are increasingly used for organizational sustainability reporting. The approach aligns with academic efforts such as Preist et al. [4], who analyzed how interface design in YouTube affects energy demand, and with the Green Software Foundation’s Software Carbon Intensity (SCI) specification, which formalizes functional units and allocation principles for shared infrastructure.
@@
Although these frameworks facilitate transparency and comparability, they rely on several simplifying assumptions. Notably, they carry over the sector-level practice of using averaged end-user device parameters by assuming constant device power, alongside assumptions of uniform user behavior and linear relationships between time, data volume, and energy use.
@@
II-C Software-Level Efficiency
An orthogonal and very active stream of research in software-level energy efficiency has produced a range of tools and methodologies for analyzing and improving the energy behavior of programs. While these cannot be used for reporting since they do not provide the absolute energy consumed, but instead provides relative energy consumed and can therefore provide insights to which programs consume more energy. A series of tools EACOF [15], PowerAPI [16], Keppler [17] –a comparison of these and other related tools are given by Jay et al. [18]– enabled applications to access power data at runtime, fostering awareness of energy use during development. These tools uses Intel’s RAPL [19, 20] which estimates the energy consumed with relative precision and we cannot use this approach for reporting.
II-D Measurement Frameworks
@@
More recently the research has also focused on aligning the methodology used when measuring, e.g., Mancebo et al. [21] and Guldner et al. [5]. Mancebo et al. have focused on the procedure and documentation of the procedure and we follow the recommendations and provide the full information for our experiment in an open repository EUD-energy-model [^2] See https://github.com/SustainableSoftware/EUD-energy-model.git. Guldner et al. propose a standardized process for measuring software resource efficiency under representative user scenarios. Their rationale is that meaningful and reproducible results require measurements grounded in realistic user interactions rather than synthetic benchmarks. This principle directly aligns with our methodological choice to evaluate energy consumption through end-to-end user flows across web services.
II-E Web-Specific Energy Studies
Prior work has examined web-service energy use through both modeling and measurement. Schien et al. [6] modeled online media use across text-based, interactive, and streaming content and found that end-user devices—particularly laptops—dominate overall energy consumption, motivating categorization by interaction type and data intensity, which we also adopt.
Subsequent studies examined factors driving variability. Suski et al. [7] showed that user choices (e.g., device and video resolution) affect streaming emissions. Methodologically, Katsenou et al. [22] compared external and RAPL-based energy measurements for video compression, and Herglotz et al. [23] developed synchronized measurement procedures for video communication, showing that RAPL correlates with external measurements but remains inaccurate. In our data, one news flow included embedded video playback, enabling comparison with these controlled observations.
@@ II-E Web-Specific Energy Studies
Collectively, prior work demonstrates that web-service energy use is heterogeneous and context dependent. However, few of these studies examine whether the simplified assumptions used in service-level reporting models can capture such variability. Our work addresses this gap by providing measured, category-based data to test the validity of those assumptions under realistic web interactions.
III Methodology
This section describes the essential parts of our experimental design, data collection, and model evaluation procedures used to address the research questions.
All anonymized user-flow descriptions, hardware, physical setup, scripts, software versions, datasets, and analysis scripts are available in the open repository EUD-energy-model [^3] See https://github.com/SustainableSoftware/EUD-energy-model.git.
Figure 1: The setup for the experiment with 4 Laptops, two Siglents and two data-collecting computers with user-controlled synchronization software.
III-A Websites and User Flows
We selected eight typical websites in four categories:
  1. C1:
@@ 2. C2:
  3. C3:
    Navigation: google.com/maps, bing.com/maps; and
  4. C4:
@@
For each category, we defined realistic user flows that mimic everyday use—searching for a product, reserving a hotel room, planning a route, or reading a news article—so that equivalent functionality was tested across services. Streaming and other parameter-dependent services were deliberately excluded to ensure comparability of interaction length and structure.
III-B Devices
@@
Experiments were performed on four Intel laptops representing different hardware generations and OEM configurations:
  1. L1:
    Dell Latitude 7490 (8th gen, UHD 620),
  2. L2:
    HP EliteBook 830 G6 (8th gen, alternate firmware),
  3. L3:
    Dell Latitude E5250 (5th gen, compact chassis), and
  4. L4:
    Dell Latitude E5550 (5th gen, larger chassis).
@@ 4. L4:
This selection captures processor-generation and design-level power variation.
III-C Participants and Experimental Design
@@
Ten participants completed the experiment; one participant (P10) contributed only two runs and was excluded from the mixed-effects analyses to ensure stable estimation of random effects. The remaining nine participants provided between eight and twenty-four valid runs each. The design crossed eight user flows with four laptops, and every flow–computer combination was executed by at least four and up to seven distinct participants, ensuring balanced coverage, see EUD-energy-model [^4] https://github.com/SustainableSoftware/EUD-energy-model.git. To counter potential learning or fatigue effects while maintaining full coverage, the order of user flows and the allocation of laptops were randomized under constraints that ensured each flow–device combination was represented across participants. Each flow was repeated three times per device, yielding a total of $164$ valid observations.
III-D Experimental Setup and Measurements
@@
All laptops ran identical operating-system images (Ubuntu Desktop 24.04.3 LTS) with default power settings and brightness. The same Chromium browser version was used in incognito mode, and caches were cleared before start. Each measurement recorded device power (W) via a Siglent SPD3303X-E programmable power supply and task duration (s), allowing energy to be computed as integrated power over time. The external power-supply measurements was user-synchronized via task start and stop events to ensure precise alignment between user interaction and power logging. The setup can be seen in Figure 1. Key sources of potential variance included machine architecture, user execution behavior, and learning effects, which were addressed through randomization and replication. Potential influences of these factors and the sample size are discussed in Section V.
III-E Model Evaluation
@@
To evaluate how accurate and precise sector-level reporting models represent actual end-user device energy, we implemented both the Digst and DIMPACT formulations. In both frameworks, energy use is estimated as constant device power multiplied by interaction duration ($E_{\mathrm{model}}=P_{\mathrm{avg}}\times t$).
@@
While Digst in general uses [1, Table 1 ] a composite averaging over laptops (22 W), tower desktops (87 W), all-in-one PCs (81 W), and monitors (31 W) as the constant device power [^5] The workload adjusted average power ratings by Urban et al. [24]. we have extracted the laptop power of 22 W and use this as comparison for our experiment. In comparison, DIMPACT have chosen the lower constant 15 W for web-browsing scenarios [2, 3] [^6] The power ratings by Urban et al. [24] without workload adjustment.. We will use these constant device powers (22W and 15W) when evaluating energy models.
IV Results
Figure 4 summarizes the measured energy used by the laptops while participants completed each user flow. Energy use varied most for shopping flows (Amazon and Ebay), whereas navigation flows (map-based tasks) had a comparatively consistent low usage.
Our results show (RQ1) systematic, duration-proportional overestimation in the Digst and DIMPACT constant-power models for realistic browser flows; and (RQ2) that adding service-category and device parameters substantially improves precision, with category-level calibration retaining most of the benefit at low complexity.
Figure 2: Measured (x-axis) versus DIMPACT’s modeled energy consumption (y-axis). The red dashed line indicates perfect prediction ($y=x$). Many points lie above the perfect prediction; DIMPACT mostly overestimates the energy consumption.
IV-A Model Accuracy and Error Growth (RQ1)